Ringers solution is used in physiology experiments. One liter of the solution contains 6 5g NaCl, 0.20g NaHCO, 0.10g CaCl2 and 0.10g KC1. Assuming that one liter of water (with a density of 1.00g ml) is used to prepare the solution: a. Calculate the mass fraction and weight percent of NaHCOs in the solution b. Calculate the mass fraction and weight percent of NaCl in the solution c. Calculate the concentrations (in ppm) of K ions and Ca ons in the solution.
Answer #1

a)

Mass fraction and weight percent of NaHCO3 in the solution.

We have 0.20g NaHCO3 in 1 L solution.

We will have to add mass of every component.

6.5g NaCl + 0.20g NaHCO3 + 0.10g CaCl2 + 0.10 g KCl + 1000 g water = 1006.9 g

Mass fraction = 0.20g/1006.9g = 0.00019

Weight percent = 0.20g/1006.9g * 100 = 0.019 %

b)

Mass fraction and weight percent of NaCl in the solution.

We have 6.5g NaCl in 1 L solution.

We will have to add mass of every component.

6.5g NaCl + 0.20g NaHCO3 + 0.10g CaCl2 + 0.10 g KCl + 1000 g water = 1006.9 g

Mass fraction = 6.5g/1006.9g = 0.00645

Weight percent = 6.5g/1006.9g * 100 = 0.645 %

c)

We have 0.10g KCl in our solution.

1) Convert to grams of KCl per L of solution.

KCl transferred is 0.10g and the solution's volume is 1 L (1000 mL), then

(0.10g)/(1 L) = 0.1g of KCl per L

2) Calculate the molarity of KCl solution.

Given MW 74.56 g/mol.

(0.1g KCl)*(1 mol KCl / 74.56 g/mol) = 0.0013 mol / L = 0.0013 M KCl solution

3) Find the concentration of solution in ppm of K.

Given KCl (aq) --> K+ (aq) + Cl- (aq)

Molarity calculated is 0.0013 mol/L for KCl.

0.0013 mol KCl * (1 mol K+ / 1 mol KCl) * (39.10 g K+ / 1 mol K+) = 0.0508 g K+ / L

In ppm 0.0508 g K+ * (1000 mg / 1 g) = 50.83 mg / L = 50.83 ppm